OpenStudy (anonymous):

Write the slope-intercept equation for the line with slope 1/2 and y-intercept (0, -5)

OpenStudy (mysesshou):

sorry, was on phone. This one?

OpenStudy (anonymous):

yes and no problem

OpenStudy (mysesshou):

Ok, so slope intercept form is Y = m*x + b m is the slope b is the y intercept So for your problem, the slope = m = 1/2 and the y intercept is b = -5 then put into the equation Y = (1/2)*x + (-5) Y=(1/2)*x - 5

OpenStudy (mysesshou):

works ok?

OpenStudy (anonymous):

so does this have an = in it or do i leave it like that

OpenStudy (mysesshou):

Y = (1/2)*x - 5 Yes, this has a = sign

OpenStudy (anonymous):

i see that i have to graph the line as well

OpenStudy (mysesshou):

ok. so plot the point (0,-5)

OpenStudy (mysesshou):

Then count up one square and to the right two squares and draw another point. connect the dots.

OpenStudy (mysesshou):

and draw a line extending beyond the dots too. you can keep counting up one, to the right two places

OpenStudy (mysesshou):



OpenStudy (anonymous):

nice drawing

OpenStudy (mysesshou):

Y = (1/2)*x - 5 if y = 0, solve for x for another point: 0 = (1/2)*x-5 5=(1/2)*x x=10 so you can plot the point (10,0) also

OpenStudy (anonymous):

so i can plot both or either or?

OpenStudy (mysesshou):